OpenStudy (anonymous):

Evaluate the integral: (integral sign) (5 sin xcos xdx)

OpenStudy (anonymous):

If you can show me how to make the integral notation, I'd be more than willing to run your questions into Maple 14.

OpenStudy (amistre64):

this is do-able :)

OpenStudy (amistre64):

use change of variable... u = sin(x) du=cos(x) dx dx = du/cos(x) 5 u cos(x) du ----------- cos(x) [S] 5u du

OpenStudy (amistre64):

5u^2 ---- + C 2 substitute the u = sin(x) 5sin(x)^2 -------- + C ; but check to see if I m right 2

OpenStudy (anonymous):

Sorry, I know nothing. I'm just a freshman in high school. What are you saying when you write "(integral sign) (5 sin xcos xdx)"

OpenStudy (amistre64):

its calculus

OpenStudy (amistre64):

it means you want an equation that will derive down to "5sin(x)cos(x)"

OpenStudy (anonymous):

Thanks guys! ^_^

OpenStudy (amistre64):

was I right? :)

OpenStudy (anonymous):

Yes you're right ^_^

OpenStudy (amistre64):

yay!!! ..... that means im smart :) now whered I leave my keys......